Calculate the money multiplier from the following information:
Currency drain ratio: 20% Reserve ratio: 10% Deposit: R60 000
1
Expert's answer
2018-09-06T15:55:08-0400
The money multiplier from the following information is: m = 1/rr = 1/0.1 = 10.